-
Find the equation of the line that contains the given point and has the given slope.
P(6,2), m=-5/6

Answer by ikleyn(52781)

.

An equation of a straight line in a coordinate plane which has the slope  m  and passes through the given point  P = (a,b)  is

        y - b = m*(x-a).                  (1)


Simply substitute the given data into equation  (1),  and you will get

    y-2 = (-5/6)*(x-6).         (2)


It is the equation in the  "slope intercept form".


If you want to get the standard form equation,  simplify  (2):

   6y-12 = -5x + 30,    or,  equivalently,    5x + 6y = 30 + 12,

or,  finally,   5x+6y = 42. 

Solved.

-------------

See the lesson in this site
    - Equation for a straight line having a given slope and passing through a given point